Wine goes bad soon after opening because the ethanol (CH3CH2OH) in it reacts with oxygen gas O2 from the air to form water H2O and acetic acid CH3COOH the main ingredient of vinegar.

What mass of oxygen gas is consumed by the reaction of 2.1g of ethanol?

 


1
Expert's answer
2021-07-26T06:33:59-0400

CH3CH2OH(g)+O2(g)"\to" H2O(l)+CH3COOH(l)

moles of ethanol=2.1/46=0.04565moles

Mass of Oxygen=1:1

Mass of Oxygen=32(0.04565)= 1.4609g